Find the solution to each system by the substitution method.
2x-5y=-1
x-3y=-3

Answer by checkley77(12844)

2x-5y=-1
x-3y=-3
x=3y-3
2(3y-3)-5y=-1
6y-6-5y=-1
y=-1+6
y=5 ans.
x=3*5-3
x=15-3
x=12 ans.
Proof:
2*12-5*5=-1
24-25=-1
-1=-1